An LRC series circuit has L = 20 henries, R = 180 ohms, C = 1/280 farad, and an applied voltage E(t) = 10 sin t. Assuming no initial charge on the capacitor, but an initial current of 1 Ampere at t = 0 when the voltage is first applied, find the steady-state charge q(t) on the capacitor and the steady-state circuit current i(t).
1
Expert's answer
2021-10-11T15:47:28-0400

"LQ''+RQ'+Q\/C=E(t)"


"20Q''+180Q'+280Q=10sint"

"k^2+9k+14=0"


"Q=\frac{-9\pm \sqrt{81-56}}{2}"


"k_1=-7,\ k_2=-2"


Homogeneous Solution:


"Q=c_1e^{-7t}+c_2e^{-2t}"


Particular Solution:


"Q=Asint+Bcost"

"Q'=Acost-Bsint"

"Q''=-Asint-Bcost"

"-Asint-Bcost+9(Acost-Bsint)+14(Asint+Bcost)=10sint"

"(-A-9B+14A)sint+(-B+9A+14B)cost=10sint"


"9A+13B=0"

"13A-9B=10"


"A=-13B\/9"

"-169B-81B=90"

"B=-9\/25"

"A=13\/25"

"Q=\frac{13}{25}sint-\frac{9}{25}cost"


General Solution:


"Q=c_1e^{-7t}+c_2e^{-2t}+\frac{13}{25}sint-\frac{9}{25}cost"


"I=\frac{dQ}{dt}=-7c_1e^{-7t}-2c_2e^{-2t}+\frac{13}{25}cost+\frac{9}{25}sint"